an electric dipole consists of two oppsite charges of magnitude 1uc separted by distance of 2cm the dipole is placed in an electric field 10to power -15 vm-1the max torque does the field exert on the dipole

Grade:12

2 Answers

Husna Begum
11 years ago

Torque = P(cross product) E

P = q*d = (10^-6)*0.02 = 2* 10^-8

E = 10^-15

torque = PEsinx = 2 * 10^-23 * 1 (since max value of sinx is 1 )

ANSWER = 2 * 10^ -23 

ashish badoni
11 years ago

torque=p.Esin(theta)

here,

p=q*distance b/w them

E=10raise to power -15

sin(theta) has to be maximum as torque should be max. it implies theta=90

now,

p=10 raise to power 6*2*10 raise to the power -2

=10 raise to power 4.

it implies torque=10^4*10^-15

=10^-11(answer).

(please ignore my long writings:-})